OpenStudy (anonymous):

When 150.0 g of ZnS are burned in excess oxygen, 68.5 g of ZnO are produced according to the equation: 2ZnS + 302 -> 2ZnO + 2SO2 What is the percent yield for this reaction?

OpenStudy (jfraser):

Use stoichimetry to find out how much ZnO "should have been" made, and compare that to the amount that was actually made.\[150.0g Zn * (\frac{1 mol Zn}{MM Zn})*(\frac{mol ZnO}{mol Zn})*(\frac{MMZnO}{mol ZnO})\] where MM is the molar mass of the stuff in that space. When you find an amount this way, you are finding a theoretical yield. Percent yield is\[%Y = \frac{actual yield}{theoretical yield}\]

OpenStudy (anonymous):

%yield is the ratio of actual yield/theoretical yield.